OpenStudy (anonymous):

marathon study sesh. woot woot. ok derive ln(5^x/x^2) i got ln5- 2/x

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

word

OpenStudy (anonymous):

\[\frac{d}{dx}\ln \frac{5^x}{x^2}\] Is that your derivative?